# Holomorphic Yang-Mills fields on $B$-branes

### v1

@@ 3.4. The homotopy category
The gauge field $\psi$ on the brane ${\mathscr{F}}^{\bullet}$ determines on each ${\mathscr{O}}$-module ${\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})$ a holomorphic connection $\vartheta^{j}$.
Definition 4.
The gauge field $\psi$ is called flat, if the curvature of the connection $\vartheta^{j}$ vanishes, for all $j$.
@@
$\xi:=\phi-\psi\in{\rm Hom}_{D^{b}(Y)}\big{(}{\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})\big{)},$
and denote by $\hat{\xi}$ the corresponding morphism ${\mathscr{F}}^{\bullet}\to\widehat{{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})}$ in the category $K^{b}(Y)_{\rm coh}.$ Thus, $\hat{\xi}$ determines a well defined morphism of ${\mathscr{O}}$-modules between the cohomologies, $\boldsymbol{\xi}^{j}:{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\to{\mathscr{H}}^{j}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})).$
On the other hand, as $\phi$ and $\psi$ are gauge fields, $\hat{\pi}\hat{\xi}=0$. Hence, $\boldsymbol{\xi}^{j}$ defines morphisms of ${\mathscr{O}}$-modules $\zeta^{j}({\xi}):{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\to\Omega^{1}({\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})).$ We denote by $\vartheta^{j}$ and $\chi^{j}$ the connections on ${\mathscr{H}}^{j}({\mathscr{F}})$ determined by $\hat{\psi}$ and $\hat{\phi}$, respectively. Since $\boldsymbol{\xi}^{j}=(\boldsymbol{\phi}^{j}-\boldsymbol{\eta}^{j})-(\boldsymbol{\psi}^{j}-\boldsymbol{\eta}^{j})$, it follows that $\chi^{j}=\vartheta^{j}+{\zeta}^{j}(\xi)$. In short, $\zeta^{j}(\xi)$ is the “variation” on the connection $\vartheta^{j}$ induced by the “variation” $\xi$ of the gauge field $\psi$.
3.5. Yang-Mills fields on a brane.
The result deduced in the following paragraph gives us a suggestion for the definition of the Yang-Mills functional over the gauge fields on a brane over a Hodge manifold.
@@
Let ${\mathscr{A}}$ be a coherent sheaf on the Hodge manifold $Y,$ and $\alpha:{\mathscr{A}}\to\Omega^{1}({\mathscr{A}})$ a holomorphic connection on ${\mathscr{A}}$. Denoting by ${\mathcal{S}}_{\mathscr{A}}$ the singular set of ${\mathscr{A}}$, on $Y\setminus{\mathcal{S}}_{\mathscr{A}}$ we define differential form
$\Phi({\mathscr{A}},\alpha):={\rm tr}\big{(}K_{\alpha}\wedge\star K_{\alpha}\big{)}\in\Gamma(Y\setminus{\mathcal{S}}_{\mathscr{A}},\,\Omega^{\rm top}\big{)},$
$K_{\alpha}$ being the curvature of $\alpha$ considered as an ${\rm End}(A)$-valued $2$-form.
@@
Proof. Let $y_{0}\in Y\setminus{\mathcal{S}}$. As the exact sequence splits locally on $Y\setminus{\mathcal{S}}$, there exists an open neighborhood $U$ of $y_{0}$ such that $g|_{U}$, in the sequence of locally free modules $0\to{\mathscr{A}}|_{U}\overset{f|_{U}}{\to}{\mathscr{B}}|_{U}\overset{g|_{U}}{\to}{\mathscr{C}}|_{U}\to 0,$ has a right inverse $h$.
Let $a$ be a frame for ${\mathscr{A}}|_{U}$, then $\alpha(a)={\sf A}\cdot a$, where ${\sf A}$ is a matrix of $1$-forms on $U$. Furthermore, $a$ can be chosen so that ${\sf A}(y_{0})=0$. Similarly, let $c$ be a frame for ${\mathscr{C}}|_{U}$, then $\gamma(c)={\sf C}\cdot c$ and we choose $c$ so that ${\sf C}(y_{0})=0$. From the splitting, it follows that $\{f(a),\,h(c)\}$ is a frame for ${\mathscr{B}}|_{U}$. By the compatibility of the connections with $f$ and $g$,
@@
Since ${\sf A}(y_{0})=0$ and ${\sf C}(y_{0})=0$, the matrix of $K_{\alpha}(y_{0})$, of the curvature of $\alpha$ at the point $y_{0}$, is $d{\sf A}$. Analogously, the matrix of $K_{\gamma}(y_{0})$ is $d{\sf C}$. The one of $K_{\beta}(y_{0})$ is the exterior derivative of (3.25), since ${\sf M}\wedge{\sf M}=0$ at $y_{0}$. Then
${\rm tr}\big{(}K_{\beta}(y_{0})\wedge\star K_{\beta}(y_{0})\big{)}$ $={\rm tr}(d{\sf A}\wedge\star d{\sf A})+{\rm tr}(d{\sf C}\wedge\star d{\sf C})$
$={\rm tr}\big{(}K_{\alpha}(y_{0})\wedge\star K_{\alpha}(y_{0})\big{)}+{\rm tr}\big{(}K_{\gamma}(y_{0})\wedge\star K_{\gamma}(y_{0})\big{)}.$
As $y_{0}$ is an arbitrary point of $Y\setminus{\mathcal{S}}$, it follows the proposition. ∎
Let $({\mathscr{G}}^{\bullet},\,{\delta}^{\bullet})$ be a bounded complex of coherent sheaves on the manifold $Y$. Let $\nabla^{\bullet}$ be a family of holomorphic connections, compatible with the operators $\delta^{\bullet}$. That is, $\nabla^{i}:{\mathscr{G}}^{i}\rightarrow\Omega^{1}({\mathscr{G}}^{i})$ is a holomorphic gauge field on the coherent sheaf ${\mathscr{G}}^{i}$ such that
$(1\otimes\delta^{i})\nabla^{i}=\nabla^{i+1}\delta^{i}.$
Hence, $\nabla^{i}({\mathscr{K}er}({\delta}^{i}))\subset{\mathscr{K}er}(1\otimes{\delta}^{i})$ and a similar relation for the image ${\mathscr{I}m}(\delta^{i-1})$. It follows that $\nabla^{i}$ induces a connection $\theta^{i}$ on the cohomology
$\theta^{i}:{\mathscr{H}}^{i}({\mathscr{G}}^{\bullet})\to\Omega^{1}({\mathscr{H}}^{i}({\mathscr{G}}^{\bullet})).$
Obviously, the restrictions of $\nabla^{i}$ determine connections on ${\mathscr{K}er}(\delta^{i})$ and ${\mathscr{I}m}(\delta^{i+1})$, respectively. One has the exact sequence
@@
Denoting with ${\mathcal{S}}$ the union of the singular sets of the sheaves ${\mathscr{G}}^{i}$, then on $Y\setminus{\mathcal{S}}$
(3.28) $\sum_{i}(-1)^{i}{\rm tr}\big{(}K_{\nabla^{i}}\wedge\star K_{\nabla^{i}}\big{)}=\sum_{i}(-1)^{i}{\rm tr}\big{(}K_{\theta^{i}}\wedge\star K_{\theta^{i}}\big{)}.$
@@
$\sum_{i}(-1)^{i}\Phi({\mathscr{G}}^{i},\,\nabla^{i})=\sum_{i}(-1)^{i}\Phi({\mathscr{H}}^{i},\,\theta^{i}).$
3.5.2. The Yang-Mills functional.
@@
    As a gauge field is a homotopy class of a morphism of complexes, it seems convenient to move on the cohomology of these complexes.
@@
    Let $E^{\bullet}$ be a bounded complex of Hermitian vector bundles over the Kähler manifold $Y,$ and $\nabla^{\bullet}$ a family of connections compatible with the Hermitian metrics and the coboundary operators. Denoting by $H^{i}(E^{\bullet})$ the cohomology bundles, there exist connections $\theta^{i}$ on those bundles, induced by the family $\nabla^{\bullet}$. By Corollary 3 together with (2.19), one has the following equality of Euler-Poincaré type.
@@
On the basis of the above considerations, it seems appropriate to define the value of the Yang-Mills functional on the gauge $\psi$ on the brane ${\mathscr{F}}^{\bullet}$ as $\sum_{i}(-1)^{i}\lVert K_{\vartheta_{i}}\rVert^{2}.$
More precisely, taking into account Proposition 11, we adopt the following definitions.
@@
The brane $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ is called a Hermitian brane, if the cohomology sheaves ${\mathscr{H}}^{j}$ are Hermitian.
Let $({\mathscr{F}}^{\bullet},\,{\delta}^{\bullet})$ be a Hermitian brane on the Hodge manifold $Y$. Given a gauge field $\psi$ on the brane $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$, by Proposition 11, one has the family of curvatures ${\mathcal{K}}_{\vartheta^{i}}$ of the connections induced on the cohomologies. Whose norms $\lVert{\mathcal{K}}_{\vartheta^{i}}\rVert$ are defined in accordance with (2.18).
@@
Given a gauge field $\psi$ on the Hermitian $B$-brane $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$, if the sheaves ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$, are reflexive, we define the value of the Yang-Mills functional at $\psi$ by
@@
Note that, if $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ is an acyclic complex, then the Yang-Mills functional for this complex is identically zero.
@@
$\vartheta^{i}=\tilde{\vartheta}^{i}+\sum_{a}\lambda_{a}\zeta^{i}_{a},$
with $\zeta^{i}_{a}\in{\rm Hom}({\mathscr{H}}^{i},\,\Omega^{1}({\mathscr{H}}^{i}))$. The corresponding curvatures satisfy
$K_{\vartheta^{i}}=K_{\tilde{\vartheta}^{i}}+\sum_{a}\lambda_{a}\tilde{\vartheta}^{i}(\zeta^{i}_{a})+\sum_{a,b}\lambda_{a}\lambda_{b}\zeta^{i}_{a}\wedge\zeta^{i}_{b}.$
@@
Let $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ be a $B$-brane, such that the ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$ are reflexive sheaves. If $\phi$ and $\psi$ are gauge fields on ${\mathscr{F}}^{\bullet}$ and $\xi=\phi-\psi$, using the notations introduced in Remark 7, the connections on the cohomologies induced by $\phi$ and $\psi$ satisfy $\chi^{j}(\xi)=\vartheta^{j}+\zeta^{j}(\xi)$, with
$\zeta^{j}(\xi)\in\Gamma(Y\setminus{\mathcal{S}},\,\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{E}nd}(H^{j})),$
@@
$\vartheta^{j}_{\epsilon}=\vartheta^{j}+\epsilon\zeta^{j}.$
@@
$K_{\vartheta^{j}_{\epsilon}}=K_{\vartheta^{j}}+\epsilon\vartheta^{j}({\zeta^{j}})+O((\epsilon)^{2}),$
$\vartheta^{j}({\zeta^{j}})$ being the covariant derivative of ${\zeta^{j}}$ considered as a section of $\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{E}nd}(H^{j})$.
@@
(3.32) $(1/2)\frac{d}{d\epsilon}\mathcal{YM}(\psi_{\epsilon})\Big{|}_{\epsilon=0}=\sum_{j}(-1)^{j}\langle K_{\vartheta^{j}},\,\vartheta^{j}(\zeta^{j})\rangle=0,$
for all any variation of $\psi$.
In particular, if $\vartheta^{i}$ es a Yang-Mills field for all $i,$ then by (2.23)
(3.33) $\langle K_{\vartheta^{i}},\,\vartheta^{i}(\zeta^{i})\rangle=0,$
and we have the following proposition.
@@
Let $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ be a Hermitian brane, such that the sheaves ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$ are reflexive, and let $\psi$ gauge field on ${\mathscr{F}}^{\bullet}$. If $\vartheta^{i}$ is a Yang-Mills field on ${\mathscr{H}}^{i}$ for all $i$, then $\psi$ is a stationary point of the Yang-Mills functional; i.e. $\psi$ is a Yang-Mills field on the brane.
The following theorem is a converse to Proposition 14.
@@
Let ${\mathscr{F}}^{\bullet}$ be a Hermitian $B$-brane as in Proposition 14. If $\nabla^{\bullet}$ is a Yang-Mills field on ${\mathscr{F}}^{\bullet},$ then the connection $\vartheta^{j}$ induced on ${\mathscr{H}}^{j}$ is a Yang-Mills field on this sheaf.
Proof. We will consider $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ as an object of the category $K^{b}_{\rm coh}(Y)$; that is, we assume that $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ is a complex of injective ${\mathscr{O}}$-modules such that its cohomology is bounded and coherent.
As ${\mathscr{K}er}(\delta^{i})$ is a submodule of the injective ${\mathscr{O}}$-module ${\mathscr{F}}^{i}$, then ${\mathscr{K}er}(\delta^{i})$ is a retract of ${\mathscr{F}}^{i}$ and thus it is also an injective ${\mathscr{O}}$-module. Therefore, the following short exact sequence of coherent ${\mathscr{O}}$-modules
@@
splits. That is, ${\mathscr{F}}^{i}\simeq{\mathscr{K}er}(\delta^{i})\oplus{\mathscr{C}oim}(\delta^{i}).$
Since ${\mathscr{I}m}({\delta}^{i-1})$ is a retract of the injective ${\mathscr{O}}$-module ${\mathscr{K}er}(\delta^{i})$, the following short exact sequence also splits
$0\to{\mathscr{I}m}({\delta}^{i-1})\to{\mathscr{K}er}(\delta^{i})\to{\mathscr{H}}^{i}\to 0$
Thus,
@@
where ${\mathscr{G}}^{i}$ is isomorphic to the direct sum of ${\mathscr{C}oim}(\delta^{i})$ and ${\mathscr{I}m}(\delta^{i-1})$. As ${\mathscr{H}}^{i}$ and ${\mathscr{G}}^{i}$ are summands in a direct sum decomposition of an injective ${\mathscr{O}}$-module, they are also injective.

### v2

@@ 3.1.2. Yang-Mills fields on reflexive sheaves
The holomorphic gauge field $\nabla$ on the Hermitian reflexive sheaf ${\mathscr{F}}$ is a Yang-Mills field iff its curvature $K_{\nabla}\in\Gamma(Y,\,\Omega^{2}\otimes_{\mathscr{O}}{\rm End}(F))$ is orthogonal to the vector space ${\rm Im}\,(^{(1)}\nabla)$.
3.1.3. The case ${\rm rank}\,{\mathscr{F}}=1$
In this case the reflexive sheaf ${\mathscr{F}}$ is a locally free sheaf [13, Prop. 1.9]. Then ${\mathscr{E}nd}({\mathscr{F}})$ is the sheaf associated to the trivial line bundle ×${\mathbb{C}}\times X\to X$. If $s$ is a local frame of the corresponding line bundle $F,$ a holomorphic connection $\nabla$ on ${\mathscr{F}}$ on this frame is determined by a ${\mathbb{C}}$-valued $1$-form $A$, $\nabla s=As$. In this frame ${}^{(p)}\nabla(\beta)=\partial\beta+A\wedge\beta-(-1)^{p}\beta\wedge A=\partial\beta,$ for any ${\mathbb{C}}$-valued $p$-form $\beta$. That is,
(3.11) $^{(p)}\nabla=\partial.$
On the other hand, the Bianchi’s identity reduces to $\partial K_{\nabla}=0$. If furthermore, $\nabla$ is a Yang-Mills field, from (3.10) it follows $\partial^{\dagger}(K_{\nabla})=0$, where $\partial^{\dagger}$ is the adjoint of $\partial$. That is, $K_{\nabla}$ is $\partial$-harmonic. As $X$ is a Kähler manifold, $K_{\nabla}$ is also $d$-harmonic. Hence, the norm of $K_{\nabla}$ minimizes the corresponding norm in its cohomology class. That is, denoting by ${\bf c}$ the cohomology class defined by $K_{\nabla},$
3.2. Yang-Mills fields on $B$-branes
In this subsection we assume that $X$ be a smooth projective variety. Then any object of $D^{b}(X)$ is isomorphic to a bounded complex consisting of locally free sheaves [24, Sect. 36.36]. Thus, we assume that the $B$-brane ${\mathscr{F}}^{\bullet}$ on $X$ is a bounded complex of locally free sheaves on $X.$ In this case, the jet sheaf ${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ is the complex of vector bundles ${\mathscr{F}}^{\bullet}\oplus\Omega^{1}({\mathscr{F}}^{\bullet})$ endowed with ${\mathscr{O}}_{X}$-module structure (2.9).
[…]
${\mathscr{H}}^{i}({\mathscr{M}}^{\bullet})\overset{(s^{i})^{-1}}{\rightarrow}{\mathscr{H}}^{i}({\mathscr{G}}^{\bullet})\overset{f^{i}}{\rightarrow}{\mathscr{H}}^{i}({\mathscr{N}}^{\bullet}).$
[…]
Thus, a gauge field on ${\mathscr{F}}^{\bullet}$, i.e. a morphism $\psi\in{\rm Hom}_{D^{b}(X)}\big{(}{\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})\big{)}$ satisfying $\pi\circ\psi={\rm id},$ determines a unique morphism
$\psi^{i}:{\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})\to{\mathscr{H}}^{i}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}))={\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})\,\tilde{\oplus}\,\Omega^{1}({\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})),$
such that the composition with the projection ${\mathscr{H}}^{i}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}))\overset{\pi^{i}}{\to}{\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$ is the identity on ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet}).$ That is, $\psi^{i}$ is a holomorphic connection on the sheaf ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet}).$
We set $\eta^{j}$ for the morphism of abelian sheaves defined by the inclusion in the direct sum $\eta^{j}:{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\rightarrow{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\oplus\Omega^{1}({\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})).$ Hence, $\pi^{j}(\psi^{j}-\eta^{j})=0$, and thus $\psi^{j}-\eta^{j}$ defines a morphism of abelian sheaves
(3.13) $\vartheta^{j}:{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\to{\Omega^{1}}({\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})),$
which, by (2.9), satisfies the Leibniz rule. That is,
Proposition 8.
The gauge field $\psi$ on the brane ${\mathscr{F}}^{\bullet}$ determines on each ${\mathscr{O}}_{X}$-module ${\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})$ a holomorphic connection $\vartheta^{j}$.
@@ 3.2. Yang-Mills fields on $B$-branes
$\xi:=\phi-\psi\in{\rm Hom}_{D^{b}(X)}\big{(}{\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})\big{)}.$
Thus, $\xi$ determines a well defined morphism of ${\mathscr{O}}_{X}$-modules between the cohomologies, $\xi^{j}:{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\to{\mathscr{H}}^{j}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})).$
We denote by $\vartheta^{j}$ and $\chi^{j}$ the connections on ${\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})$ determined by $\psi$ and $\phi$, respectively. Since $\xi^{j}=(\phi^{j}-\eta^{j})-(\psi^{j}-\eta^{j})$, it follows that $\xi^{j}=\vartheta^{j}-\chi^{j}$. In short, $\xi^{j}$ is the “variation” on the connection $\vartheta^{j}$ induced by the “variation” $\xi$ of the gauge field $\psi$.
The result deduced in the following paragraph gives us a suggestion for the definition of the Yang-Mills functional on the gauge fields on a brane over a Kähler manifold.
@@
Let ${\mathscr{A}}$ be a coherent sheaf on the Kähler manifold $X,$ and $\alpha:{\mathscr{A}}\to\Omega^{1}({\mathscr{A}})$ a holomorphic connection on ${\mathscr{A}}$. Denoting by ${\mathcal{S}}_{\mathscr{A}}$ the singular set of ${\mathscr{A}}$, on $Y\setminus{\mathcal{S}}_{\mathscr{A}}$ we define differential form
$\Phi({\mathscr{A}},\alpha):={\rm tr}\big{(}K_{\alpha}\wedge\star K_{\alpha}\big{)}\in\Gamma(X\setminus{\mathcal{S}}_{\mathscr{A}},\,\Omega^{\rm top}\big{)},$
where $K_{\alpha}$ is the curvature of $\alpha$, considered as an ${\rm End}(A)$-valued $2$-form.
@@
Proof. Let $x_{0}\in X\setminus{\mathcal{S}}$. As the exact sequence splits locally on $X\setminus{\mathcal{S}}$, there exists an open neighborhood $U$ of $x_{0}$ such that $g|_{U}$, in the sequence of locally free modules $0\to{\mathscr{A}}|_{U}\overset{f|_{U}}{\to}{\mathscr{B}}|_{U}\overset{g|_{U}}{\to}{\mathscr{C}}|_{U}\to 0,$ has a right inverse $h$.
Let $a$ be a frame for ${\mathscr{A}}|_{U}$, then $\alpha(a)={\sf A}\cdot a$, where ${\sf A}$ is a matrix of $1$-forms on $U$. Furthermore, $a$ can be chosen so that ${\sf A}(x_{0})=0$. Similarly, let $c$ be a frame for ${\mathscr{C}}|_{U}$, then $\gamma(c)={\sf C}\cdot c$ and we choose $c$ so that ${\sf C}(x_{0})=0$. From the splitting, it follows that $\{f(a),\,h(c)\}$ is a frame for ${\mathscr{B}}|_{U}$. By the compatibility of the connections with $f$ and $g$,
@@
Since ${\sf A}(x_{0})=0$ and ${\sf C}(x_{0})=0$, the matrix of $K_{\alpha}(x_{0})$, of the curvature of $\alpha$ at the point $x_{0}$, is $d{\sf A}$. Analogously, the matrix of $K_{\gamma}(x_{0})$ is $d{\sf C}$. The one of $K_{\beta}(x_{0})$ is the exterior derivative of (3.14), since ${\sf M}\wedge{\sf M}=0$ at $x_{0}$. Then
${\rm tr}\big{(}K_{\beta}(x_{0})\wedge\star K_{\beta}(x_{0})\big{)}$ $={\rm tr}(d{\sf A}\wedge\star d{\sf A})+{\rm tr}(d{\sf C}\wedge\star d{\sf C})$
$={\rm tr}\big{(}K_{\alpha}(x_{0})\wedge\star K_{\alpha}(x_{0})\big{)}+{\rm tr}\big{(}K_{\gamma}(x_{0})\wedge\star K_{\gamma}(x_{0})\big{)}.$
As $x_{0}$ is an arbitrary point of $X\setminus{\mathcal{S}}$, it follows the proposition. ∎
Let $({\mathscr{G}}^{\bullet},\,{\delta}^{\bullet})$ be a bounded complex of coherent sheaves on the manifold $X$. Let $\nabla^{\bullet}$ be a family of holomorphic connections, compatible with the operators $\delta^{\bullet}$. That is, $\nabla^{i}:{\mathscr{G}}^{i}\rightarrow\Omega^{1}({\mathscr{G}}^{i})$ is a holomorphic connection on the coherent sheaf ${\mathscr{G}}^{i}$ such that $(1\otimes\delta^{i})\nabla^{i}=\nabla^{i+1}\delta^{i}.$ Hence, $\nabla^{i}({\mathscr{K}er}({\delta}^{i}))\subset{\mathscr{K}er}(1\otimes{\delta}^{i})$ and a similar relation for the image ${\mathscr{I}m}(\delta^{i-1})$. It follows that $\nabla^{i}$ induces a connection $\theta^{i}$ on the cohomology $\theta^{i}:{\mathscr{H}}^{i}({\mathscr{G}}^{\bullet})\to\Omega^{1}({\mathscr{H}}^{i}({\mathscr{G}}^{\bullet})).$ Obviously, the restrictions of $\nabla^{i}$ determine connections on ${\mathscr{K}er}(\delta^{i})$ and ${\mathscr{I}m}(\delta^{i+1})$, respectively. One has the exact sequence
@@
Denoting with ${\mathcal{S}}$ the union of the singular sets of the sheaves ${\mathscr{G}}^{i}$, then $\,\sum_{i}(-1)^{i}{\rm tr}\big{(}K_{\nabla^{i}}\wedge\star K_{\nabla^{i}}\big{)}=\sum_{i}(-1)^{i}{\rm tr}\big{(}K_{\theta^{i}}\wedge\star K_{\theta^{i}}\big{)}$ on $X\setminus{\mathcal{S}}.$
@@
$\sum_{i}(-1)^{i}\Phi({\mathscr{G}}^{i},\,\nabla^{i})=\sum_{i}(-1)^{i}\Phi({\mathscr{H}}^{i},\,\theta^{i}).\qed$
3.2.2. The Yang-Mills functional
@@
    As a gauge field $\psi$ on ${\mathscr{F}}^{\bullet}$ is a class of “roofs” [9, p. 148] from the complex ${\mathscr{F}}^{\bullet}$ to ${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}),$ and equivalent roofs determine the same morphisms between the cohomologies, it seems convenient to move on the cohomology of these complexes.
@@
    Let $E^{\bullet}$ be a bounded complex of Hermitian vector bundles over the Kähler manifold $Y,$ and $\nabla^{\bullet}$ a family of connections compatible with the Hermitian metrics and the coboundary operators. Denoting by ${\mathscr{H}}^{i}(E^{\bullet})$ the cohomology sheaves, there exist connections $\theta^{i}$ on those sheaves, induced by the family $\nabla^{\bullet}$. By Corollary 3 together with (3.5), one has the following equality of Euler-Poincaré type.
@@
On the basis of the above considerations, it seems appropriate to define the value of the Yang-Mills functional on the gauge $\psi$ on the brane ${\mathscr{F}}^{\bullet}$ as $\sum_{i}(-1)^{i}\lVert K_{\vartheta_{i}}\rVert^{2}.$ More precisely, taking into account Proposition 8, we adopt the following definitions.
@@
The brane $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ is called a Hermitian brane, if the cohomology sheaves ${\mathscr{H}}^{j}$ are Hermitian ${\mathscr{O}}_{X}$-modules.
Let $({\mathscr{F}}^{\bullet},\,{\delta}^{\bullet})$ be a Hermitian brane on the Kähler manifold $X$. Given a gauge field $\psi$ on the brane $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$, by Proposition 8, one has the family of curvatures ${\mathcal{K}}_{\vartheta^{i}}$ of the connections induced on the cohomologies, whose norms $\lVert{\mathcal{K}}_{\vartheta^{i}}\rVert$ are defined in accordance with (3.4).
@@
Given a gauge field $\psi$ on the Hermitian $B$-brane $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$, if the sheaves ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$ are reflexive, we define the value of the Yang-Mills functional at $\psi$ by
@@
Note that if $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ is an acyclic complex, then the Yang-Mills functional for this complex is identically zero.
@@
$\vartheta^{i}=\tilde{\vartheta}^{i}+\sum_{a}\lambda_{a}\xi^{i}_{a},$
with $\xi^{i}_{a}\in{\rm Hom}({\mathscr{H}}^{i},\,\Omega^{1}({\mathscr{H}}^{i}))$. The corresponding curvatures satisfy
$K_{\vartheta^{i}}=K_{\tilde{\vartheta}^{i}}+\sum_{a}\lambda_{a}\tilde{\vartheta}^{i}(\xi^{i}_{a})+\sum_{a,b}\lambda_{a}\lambda_{b}\xi^{i}_{a}\wedge\xi^{i}_{b}.$
@@
Let $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ be a $B$-brane, such that the ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$ are reflexive sheaves. If $\phi$ and $\psi$ are gauge fields on ${\mathscr{F}}^{\bullet}$ and $\xi=\phi-\psi$, using the notations introduced in Remark 6, the connections on the cohomologies induced by $\phi$ and $\psi$ satisfy $\chi^{j}(\xi)=\vartheta^{j}+\xi^{j}$, with
$\xi^{j}\in\Gamma(Y\setminus{\mathcal{S}},\,\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{E}nd}(H^{j})),$
@@
$\vartheta^{j}_{\epsilon}=\vartheta^{j}+\epsilon\xi^{j}.$
@@
$K_{\vartheta^{j}_{\epsilon}}=K_{\vartheta^{j}}+\epsilon\vartheta^{j}({\xi^{j}})+O(\epsilon^{2}),$
$\vartheta^{j}({\xi^{j}})$ being the covariant derivative of ${\xi^{j}}$ considered as a section of $\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{E}nd}(H^{j})$.
@@
(3.20) $(1/2)\frac{d}{d\epsilon}\mathcal{YM}(\psi_{\epsilon})\Big{|}_{\epsilon=0}=\sum_{j}(-1)^{j}\langle K_{\vartheta^{j}},\,\vartheta^{j}(\xi^{j})\rangle=0,$
for all any variation of $\psi$. In particular, if $\vartheta^{i}$ es a Yang-Mills field for all $i,$ then by (3.10) $\langle K_{\vartheta^{i}},\,\vartheta^{i}(\xi^{i})\rangle=0,$ and we have the following proposition.
@@
Let $\psi$ gauge field on the brane $({\mathscr{F}}^{\bullet},\,\delta^{\bullet}).$ If $\vartheta^{i}$ is a Yang-Mills field on ${\mathscr{H}}^{i}$ for all $i$, then $\psi$ is a Yang-Mills field on the brane.
The following theorem is a partial converse to Proposition 11.
@@
If each ${\mathscr{F}}^{i}$ is a semisimple object in the category of coherent sheaves on $X$ and $\nabla^{\bullet}$ is a Yang-Mills field on ${\mathscr{F}}^{\bullet},$ then the connection $\vartheta^{j}$ induced on ${\mathscr{H}}^{j}$ is a Yang-Mills field on this sheaf.
Proof. Since ${\mathscr{F}}^{i}$ is semisimple the following short exact sequence of coherent ${\mathscr{O}}_{X}$-modules
@@
splits. That is, ${\mathscr{F}}^{i}\simeq{\mathscr{K}er}(\delta^{i})\oplus{\mathscr{C}oim}(\delta^{i}).$ Since ${\mathscr{K}er}(\delta^{i})$ is semisimple the exact sequence $0\to{\mathscr{I}m}({\delta}^{i-1})\to{\mathscr{K}er}(\delta^{i})\to{\mathscr{H}}^{i}\to 0$ also splits Thus,
@@
where ${\mathscr{G}}^{i}$ is isomorphic to the direct sum of ${\mathscr{C}oim}(\delta^{i})$ and ${\mathscr{I}m}(\delta^{i-1})$.